# AlignFormer: Modality Matching Can Achieve Better Zero-shot Instruction-Following Speech-LLM

### v1

@@ Abstract
Integrating speech into LLM (speech-LLM) has gaining increased attention recently. The mainstream solution is to connect a well-trained speech encoder and LLM with a neural adapter. However, the length mismatch between the speech and text sequences are not well handled, leading to imperfect modality matching between the speech and text. In this work, we propose a novel neural adapter, AlignFormer, to reduce the length gap between the two modalities. AlignFormer consists of CTC and dynamic-window QFormer layers, where the CTC alignment provides the dynamic window information for qformer layers. The LLM backbone is frozen in training to preserve its text capability, especially the instruction following capability. When training with only the ASR data, the proposed AlignFormer unlocks the instruction following capability for speech-LLM and the model can perform zero-shot speech translation (ST) and speech question answering (SQA) tasks. In fact, speech-LLM with AlignFormer can theoretically perform any tasks that the LLM backbone can deal with in the speech version. To evaluate the effectiveness of the instruction-following speech-LLM, we propose to use instruction following rate (IFR) and offer a systematic perspective for the IFR evaluation. In addition, we find that the audio position in training would affect the instruction following capability of speech-LLM and conduct an in-depth study on it. Our findings show that audio-first training achieves higher IFR than instruction-first training. The AlignFormer can achieve a near 100% IFR with audio-first training and game-changing improvements from zero to non-zero IFR on some evaluation data with instruction-first training. We believe that this study is a big step towards the perfect speech and text modality matching in the LLM embedding space.
@@ I Introduction
The most common way to achieve speech-LLM is to integrate a speech encoder into LLM with a neural adapter. The neural adapter can be either linear, convolution or transformer layers to map the speech encoder outputs to text embedding space of the LLM. For example, the authors in [13, 22] explore such architectures for a single automatic speech recognition (ASR) task and show competitive performance compared to the expert ASR model. Later on, several works, such as Qwen2-audio [15], SALMONN [14], SpeechVerse [19], and WavLLM [23], scale the number of tasks through instruction tuning. The instruction data is created by formulating various speech and audio centric tasks into a generative paradigm. Such instruction-tuned speech-LLM can do well on the seen tasks in training, while the performance on the out-of-domain/unseen tasks can be terribly bad. The model is sometimes not able to follow the task instructions not seen in the training. In addition, the above-mentioned methods will tune the LLM parameters during the instruction tuning stage for better performance on different tasks, which inevitably destroys the text capability of the LLM. It is interesting to explore how to unlock the speech understanding capabilities of LLM without hurting its text performance and this work falls into the category.
@@
As we mentioned, the LLM after post-training shows strong instruction-following capability. Freezing the LLM parameters can be an efficient way to preserve the instruction following capability of the LLM. In fact, there are several works already studying the instruction-following speech-LLM without tuning LLM. In AudioChatLLama [24], the LLM response of the audio transcription is used to train speech encoder and adapter, and the model is able to follow audio instructions. However, it is unknown whether the model can follow text instructions and whether it achieves good enough performance on speech understanding tasks. In BLSP [25], the authors propose behavior alignment to train speech-LLM with the LLM responses to the continual writing/repeat prompts of the ASR transcriptions. The behavior alignment can provide speech-LLM zero-shot capabilities for speech and language understanding tasks, but the performance of these tasks are not as competitive as the instruction-tuned model. [26] and  [27] further propose to add emotion and para-lingual information when generating LLM response so that the para-lingual information can be preserved in speech encoder. Overall, the prior methods unlock the instruction-following speech-LLM through data efforts and do not consider a well-aligned speech and text joint space. Consequently, the model performance would be limited. Moreover, we are lacking a systematic way to track the instruction following capability of the speech-LLM, and thus it is hard to evaluate the effectiveness of the instruction-following speech-LLM, especially when the model fails to follow the instructions.
@@
In this work, to evaluate the instruction-following capability of the speech-LLM, we propose to use instruction following rate (IFR) as an additional metric to evaluate the effectiveness of the instruction-following speech-LLM and offer a systematic view of calculating IFR. We study the zero-shot instruction-following speech-LLM with the Phi3.1-mini-instruct model [28]. Interestingly, we find that audio-first inputs (audio embedding + text instruction) always achieves higher IFR than instruction-first inputs (text instruction + audio embedding). This is mainly due to the reason that the Phi3 model is post-trained with instruction-first data so that audio-first training can be regarded as learning from various prompts, preventing the overfitting to the single prompt in the prefix. More importantly, we propose a novel neural adapter, AlignFormer, between the speech encoder and LLM for better speech and text modality matching. AlignFormer consists of CTC [29] and dynamic-window QFormer [14] layers to carefully handle the length mismatch between the speech and text sequences. Concretely, the CTC layer would provide the length and alignment information between the speech and text sequence [30]. The dynamic-window qformer layers (qformer with dynamic windows) would leverage the information to generate speech outputs with length and embedding space to be similar to that of the text sequence. We show for the first time the zero-shot instruction-following speech-LLM with the ASR training data only (without LLM response data). On the evaluated translation and QA tasks, AlignFormer achieves almost 100% IFR in the audio-first setting, and improves the averaged IFR from 15% to 46% in the instruction-first setting, especially the game-changing improvements from zero to non-zero IFR on Gaoka and Speaker Verification QA tasks. The emerging zero-shot capability indicates that the proposed method is possible to align the speech into the LLM embedding space, which is a big step towards the perfect modality matching between the speech and text.
@@
  • We are the first to propose using instruction following rate (IFR) to evaluate the effectiveness of instruction-following speech-LLM. The usage of IFR can provide insights into the development of instruction-following speech-LLM.
@@ III Methodologies
In this section, we introduce the methodologies to achieve zero-shot instruction-following speech-LLM. We first introduce the basic architecture of the speech-LLM and different audio positions for LLM inputs. The proposed AlignFormer is then introduced in details. The section ends with how we create the evaluation data for calculating the metric of instruction following rate (IFR).
@@ III-B AlignFormer: Speech and Text Modality Matching
III-C Instruction Following Rate: Metric and Evaluation Paradigm
When studying the instruction-following speech-LLM, we find that the instruction following capability is not stable, aka. some instructions are followed while others are not. It would be hard to track the effectiveness of the instruction-following speech-LLM with only the task specific metrics, such as word error rate (WER) for ASR, BLUE scores for speech translation (ST) and accuracy for speech question answering (SQA) tasks. The low scores may come from the model not following the instructions. We, therefore, propose to use instruction following rate (IFR) to track the instruction following capability of a speech-LLM model. The IFR is simple to calculate as the ratio of number of followed examples to the total examples in the evaluated dataset.
The difficulty of calculating IFR is how to detect if the answer is followed by the request in the prompt. To solve the problem, we attempt to create different prompts for each task and explicitly instruct the model answering in a specific format. Taking the emotion recognition task as an example, we provide the emotion categories in the prompt and inform the model to response with one of the options in a specific format. We detect the format in the answer to compute the IFR and accuracy for the emotion recognition task. Most of the close-ended SQA tasks can be evaluated in such a manner. For speech translation tasks, we detect the target language to see if the model responses in the correct language. We don’t evaluate IFR for the open-ended SQA and ASR tasks because it is hard to detect if the model follow these instructions. The prompts used for different tasks are listed in Table LABEL:tab:prompts.
@@ IV Experimental Settings
In this section, we introduce the experimental settings of this work, including the data, model, training and evaluation settings.
@@ IV-A1 Data and Model
We use the 38k hours annoymized in-house English ASR data to train the speech-LLM. For audio-first training, the task prompt is “Transcribe the audio clip into text.”. For instruction-first training, we use “Repeat exactly what the user says word by word.” as the task prompt. The reason we use this task prompt is that the LLM backbone (Phi) can follow the instruction well, where near 90% of the LLM response to the ASR transcription is with WER lower than 5%. We use the internal zero-shot TTS system to convert the text instruction to audio instruction when necessary.
We use the Phi3.1-mini-instruct [28] model [^1] https://huggingface.co/microsoft/Phi-3-mini-128k-instruct as the LLM backbone in this study. Phi3.1 model is a 3.8 billion small language model, but can achieve comparable performance to 7B models like LLaMA [3]. The speech encoder consists 3 convolutions layers and 24 confomer blocks [41] with 1024 attention dimension, 1536 feed-forward dimension, and 16 attention heads. The convolution layers contribute to a sub-sampling rate of 8, and thus 80ms token rate for adapter. Architecture-wise, we train two baseline models. The first baseline is a 2-layer MLP as the adapter without further downsample. The token rate is 80ms. The second baseline is incorporating additional window qformer layer between speech encoder and the MLP with a window size of 4. Each 4 speech encoder outputs will be further grouped into one embeeding, leading to 320ms token rate. This is a more fair comparison to AlignFormer as its token rate is about 320ms. For AlignFormer, it is 2-layer cross-attention block with one learned query. Either greedy-alignment, force-alignment, or mixed alignment is used in training. The CTC alignment is converted to attention mask functioned as the dynamic window size. The CTC task ratio is 0.3. The model parameters of speech encoder and 2-layer MLP is 450M and it is 500M for speech encoder plus AlignFormer.
@@ V-A The Effect of Audio Positions for Unlocking Instruction-following Speech-LLM
The results of IFR are reported in Table II. As shown in the table, audio-first training (E1) achieves much higher IFR compared to the instruction-first training (E2). The instruction-first training can barely follow the majority of the instructions. All the failed cases as we observe are actually doing speech recognition. Since the Phi instruct model is trained with instruction-first data, we hypothesize the reason to be that the instruction-first speech-LLM training will overfit to the ASR instruction used in training. However, audio-first training is equivalent to training with large amounts of different instructions and thus the LLM instruction-following capability is preserved.
To verify our hypotheses, we conduct the experiments of E3 and E4 to convert the text instructions into their audio version and increase the instruction diversities by using additional 4 different ASR prompts. The results in Table II show that by incorporating the audio ASR instruction, the averaged IFR increases from 15% to 34% and 46% respectively for 1 and 5 audio instructions, showing the effectiveness of audio-first training for better instruction-following speech-LLM. However, the model still cannot follow on Gaokao QA data, indicating better speech and text modality matching alignment is required.
@@ V-B AlignFormer: Achieving Better Speech and Text Modality Matching
TABLE III: The zero-shot evaluations on WavLLM data. Task metrics and IFR (%) are reported. We do not compute IFR for ASR tasks because it is hard to do so. The IFR are averaged on ST and SQA tasks. The QFormer is with 320ms token rate (window size of 4) which is close to that of AlignFormer. The token rate for E1 and E2 is 80ms. Note that the results in WavLLM paper (the second row) is with instruction tuning using in-domain training data while our results are with the zero-shot evaluation (only using ASR training data).
@@
For the ASR tasks, we use repeat prompt in the evaluation. The poor performance of E1 on Librispeech is due to the hallucinations in addition to the ASR transcriptions. In other words, the model cannot follow the exact the repeat prompt. Hence, the ASR evaluations on audio-first training are still in a zero-shot manner and E1 should have a low instruction following rate. For ST and SQA tasks, we can easily calculate the IFR using the method in Section III-C. We can observe from the table that AlignFormer can greatly improve the IFR for these tasks. For audio-first training, AlignFormer achieves a near 100%. For instruction-first training, AlignFormer improves the IFR to an average of 49%. We should specifically notice that the IFR on Gaokao QA has been improved to a non-zero IFR which is a game-changing improvement. This can be an important step towards the perfect speech and text modality matching in the LLM space and we show that it might be possible to achieve. Quality-wise, the performance of AlignFormer still encounters information loss problem because of lower token rate. For example, on CoVoST2 en-de evaluation, the BLEU scores are lower than the baselines even the IFR is higher. We still see the AlignFormer is promising because on MuST-C, the AlignFormer achieves a better BLEU than the cascades ASR + LLM system.
@@
We also have interesting findings on the SQA tasks. For the emotion tasks, there are four categories. The accuracy of zero-shot emotion evaluation shows that the model can infer the emotion from the ASR transcriptions to some degree (better than random guess with 25% for 4 categories). However, for the speaker verification task, the accuracy is similar to random guess (yes or no), indicating that the model is hard to infer speaker information from text. It provides us insights into the model development for emotion and speaker verification tasks.

### v2

@@ Abstract
Integrating speech into LLM (speech-LLM) has gaining increased attention recently. The mainstream solution is to connect a well-trained speech encoder and LLM with a neural adapter. However, the length mismatch between the speech and text sequences are not well handled, leading to imperfect modality matching between the speech and text. In this work, we propose a novel neural adapter, AlignFormer, to reduce the length gap between the two modalities. AlignFormer consists of CTC and dynamic-window QFormer layers, where the CTC alignment provides the dynamic window information for QFormer. The LLM backbone is frozen in training to preserve its text capability, especially the instruction following capability. When training with ASR data only, the proposed AlignFormer unlocks the instruction following capability for speech-LLM and the model can perform zero-shot speech translation (ST) and speech question answering (SQA) tasks. In fact, speech-LLM with AlignFormer can theoretically perform any tasks that the LLM backbone can deal with in the speech version. To evaluate the effectiveness of the instruction-following speech-LLM, we propose to use instruction following rate (IFR) and offer a systematic perspective for the IFR evaluation. In addition, we find that the audio position in training would affect the instruction following capability of speech-LLM and conduct an in-depth study on it. Our findings show that audio-first training achieves higher IFR than instruction-first training. The AlignFormer can achieve a near 100% IFR with audio-first training and game-changing improvements from zero to non-zero IFR on some evaluation data with instruction-first training. We believe that this study is a big step towards the perfect speech and text modality matching in the LLM embedding space.
@@ I Introduction
The most common way to achieve speech-LLM is to integrate a speech encoder into LLM with a neural adapter. The neural adapter can be either linear, convolution or transformer layers to map the speech encoder outputs to text embedding space of the LLM. For example, the authors in [13, 22] explore such architectures for a single automatic speech recognition (ASR) task and show competitive performance compared to the expert ASR model. Later on, several works, such as Qwen2-audio [15], SALMONN [14], SpeechVerse [19], and WavLLM [23], scale the number of tasks through instruction tuning. The instruction data is created by formulating various speech and audio centric tasks into a generative manner. Such instruction-tuned speech-LLM can do well on the seen tasks in training, while the performance on the out-of-domain/unseen tasks can be terribly bad. The model is sometimes not able to follow the task instructions not seen in the training. In addition, the above-mentioned methods will tune the LLM parameters during the instruction tuning stage for better performance on different tasks, which inevitably destroys the text capability of the LLM. It is interesting to explore how to unlock the speech understanding capabilities of LLM without hurting its text performance and this work falls into the category.
@@
As we mentioned, the LLM after post-training shows strong instruction-following capability. Freezing the LLM parameters can be an efficient way to preserve the instruction following capability of the LLM. In fact, there are several works already studying the instruction-following speech-LLM without tuning LLM. In AudioChatLLama [24], the LLM response of the audio transcription is used to train speech encoder and adapter, and the model is able to follow audio instructions. However, it is unknown whether the model can follow text instructions and whether it achieves good enough performance on speech understanding tasks. In BLSP [25], the authors propose behavior alignment to train speech-LLM with the LLM responses to the continual writing/repeat prompts of the ASR transcriptions. The behavior alignment can provide speech-LLM zero-shot capabilities for speech and language understanding tasks, but the performance of these tasks are not as competitive as the instruction-tuned model. [26] and  [27] further propose to add emotion and para-lingual information when generating LLM response so that the para-lingual information can be preserved in speech encoder. [28] reveals the similar zero-shot behaviour with ASR training data only. Overall, the prior methods unlock the instruction-following speech-LLM through data efforts and do not consider a well-aligned speech and text joint space except [28]. Consequently, the model performance would be limited. Moreover, we are lacking a systematic way to track the instruction following capability of the speech-LLM, and thus it is hard to evaluate the effectiveness of the instruction-following speech-LLM, especially when the model fails to follow the instructions.
@@
In this work, to evaluate the instruction-following capability of the speech-LLM, we propose to use instruction following rate (IFR) as an additional metric to evaluate the effectiveness of the instruction-following speech-LLM and offer a systematic view of calculating IFR. We study the zero-shot instruction-following speech-LLM with the Phi3.1-mini-instruct model [29, 30]. Interestingly, we find that audio-first inputs (audio embedding + text instruction) always achieves higher IFR than instruction-first inputs (text instruction + audio embedding). This is mainly due to the reason that the Phi3 model is post-trained with instruction-first data so that audio-first training can be regarded as learning from various prompts, preventing the overfitting to the single prompt in the prefix. More importantly, we propose a novel neural adapter, AlignFormer, between the speech encoder and LLM for better speech and text modality matching. AlignFormer consists of CTC [31] and dynamic-window QFormer [14] layers to carefully handle the length mismatch between the speech and text sequences. Concretely, the CTC layer would provide the length and alignment information between the speech and text sequence [32, 33]. The dynamic-window qformer layers would leverage the information to generate speech outputs with length and embedding space to be similar to that of the text sequence. We show for the first time the zero-shot instruction-following speech-LLM with the ASR training data only (without LLM response data). On the evaluated translation and QA tasks, AlignFormer achieves almost 100% IFR in the audio-first setting, and improves the averaged IFR from 15% to 46% in the instruction-first setting, especially the game-changing improvements from zero to non-zero IFR on Gaoka and Speaker Verification QA tasks. The emerging zero-shot capability indicates that the proposed method is possible to align the speech into the LLM embedding space, which is a big step towards the perfect modality matching between the speech and text.
@@
  • We are the first to propose using IFR to evaluate the effectiveness of instruction-following speech-LLM. The usage of IFR can provide insights into the development of instruction-following speech-LLM.
@@ III Methodologies
In this section, we introduce the methodologies to achieve zero-shot instruction-following speech-LLM. We first introduce the basic architecture of the speech-LLM and different audio positions for LLM inputs. The proposed AlignFormer is then introduced in details. The section ends with how we create the evaluation data for calculating the metric of IFR.
@@ III-B AlignFormer: Speech and Text Modality Matching
III-C IFR: Metric and Evaluation Paradigm
When studying the instruction-following speech-LLM, we find that the instruction following capability is not stable, aka. some instructions are followed while others are not. It would be hard to track the effectiveness of the instruction-following speech-LLM with only the task specific metrics, such as word error rate (WER) for ASR, BLUE scores for speech translation (ST) and accuracy for speech question answering (SQA) tasks. The low scores may come from the model not following the instructions. We, therefore, propose to use IFR to track the instruction following capability of a speech-LLM model. The IFR is simple to calculate as the ratio of number of followed examples to the total examples in the evaluated dataset.
The difficulty of calculating IFR is how to detect if the answer is followed by the request in the prompt. To solve the problem, we attempt to create different prompts for each task and explicitly instruct the model answering in a specific format. Taking the emotion recognition task as an example, we provide the emotion categories in the prompt and inform the model to response with one of the options in a specific format. We detect the format in the answer to compute the IFR and accuracy for the emotion recognition task. Most of the close-ended SQA tasks can be evaluated in such a manner. For ST tasks, we detect the target language to see if the model responses in the correct language. We don’t evaluate IFR for the open-ended SQA and ASR tasks because it is hard to detect if the model follow these instructions. The prompts used for different tasks are listed in Table LABEL:tab:prompts.
@@ IV Experimental Settings
In this section, we introduce the experimental settings of this work, including the data, model, training and evaluation.
@@ IV-A1 Data and Model
We use the 38k hours anonymized in-house English ASR data to train the speech-LLM. For audio-first training, the task prompt is “Transcribe the audio clip into text.”. For instruction-first training, we use “Repeat exactly what the user says word by word.” as the task prompt. The reason we use this task prompt is that the LLM backbone (Phi) can follow the instruction well, where near 90% of the LLM response to the ASR transcription is with WER lower than 5%. We use the internal zero-shot TTS system to convert the text instruction to audio instruction when necessary.
We use the Phi3.1-mini-instruct [29] model [^1] https://huggingface.co/microsoft/Phi-3-mini-128k-instruct as the LLM backbone in this study. Phi3.1 model is a 3.8 billion small language model, but can achieve comparable performance to 7B models like LLaMA [3]. The speech encoder consists 3 convolutions layers and 24 conformer blocks [44] with 1024 attention dimension, 1536 feed-forward dimension, and 16 attention heads. The convolution layers contribute to a sub-sampling rate of 8, and thus 80ms token rate for adapter. Architecture-wise, we train two baseline models. The first baseline is a 2-layer MLP as the adapter without further downsample. The token rate is 80ms. The second baseline is incorporating additional window qformer layer between speech encoder and the MLP with a window size of 4. Each 4 speech encoder outputs will be further grouped into one embeeding, leading to 320ms token rate. This is a more fair comparison to AlignFormer as its token rate is about 320ms. For AlignFormer, it is 2-layer cross-attention block with one learned query. Either greedy-alignment, force-alignment, or mixed alignment is used in training. The CTC alignment is converted to attention mask functioned as the dynamic window size. The CTC task ratio is 0.3. The model parameters of speech encoder and 2-layer MLP is 450M and it is 500M for speech encoder plus AlignFormer.
@@ IV-B2 Evaluation Details
Note that the IFR metric should be very sensitive to prompts. However, once the prompt and detection criteria are determined, the IFR results are stable and reliable. In this work, we only evaluate IFR with the prompt shown in Table LABEL:tab:prompts. It is interesting to evaluate with more prompt variations for general conclusions. However, creating a prompt pool that can both easily detect the following behavior is not easy. We leave it for future work.
@@ V-A The Effect of Audio Positions for Unlocking Instruction-following Speech-LLM
The results of IFR are reported in Table II. As shown in the table, audio-first training (E1) achieves much higher IFR compared to the instruction-first training (E2). The instruction-first training can barely follow the majority of the instructions. All the failed cases as we observe are performing ASR. Since the Phi instruct model is trained with instruction-first data, we hypothesize the reason to be that the instruction-first speech-LLM training will overfit to the ASR instruction used in training. However, audio-first training is equivalent to training with large amounts of different instructions and thus the LLM instruction-following capability is preserved.
To verify our hypotheses, we conduct the experiments of E3 and E4 to convert the text instructions into their audio version and increase the instruction diversities with 4 additional ASR prompts. The results in Table II show that by incorporating the audio ASR instruction, the averaged IFR increases from 15% to 34% and 46% respectively for 1 and 5 audio instructions, showing the effectiveness of audio-first training for better instruction-following speech-LLM. However, the model still cannot follow on Gaokao QA data, indicating better speech and text modality matching alignment is required.
Box 1. A Hallucination Example with Repeat Prompt in ASR Evaluation Audio_id: 2609-169640-0017.flac Reference: the john behaved beautifully and came round like a top Hypothesis: The phrase ”The John behaved beautifully. He came around like a top.” is to be repeated exactly as provided, without any alterations. The text following the phrase is cut off, but the instruction is clear that the repetition should continue from the given text. ## Response: The John behaved beautifully. He came around like a top.
@@ V-B AlignFormer: Achieving Better Speech and Text Modality Matching
TABLE III: The zero-shot evaluations on WavLLM data. Task metrics and IFR (%) are reported. We do not compute IFR for ASR tasks because it is hard to do so. The IFR are averaged on ST and SQA tasks. The QFormer is with 320ms token rate (window size of 4) which is close to that of AlignFormer. The token rate for E1 and E2 is 80ms. Note that the results in WavLLM paper (the second row) is with instruction tuning using in-domain instruction-following data so that it achieves 100% IFR, while our results are with the zero-shot evaluation (only using ASR training data). CTC out + LLM indicates that the ASR greedy search results from CTC head is fed into LLM as a fair comparison to the proposed AlignFormer, which operates in the latent space. SFT-LLM-Freeze stands for freezing LLM during supervised finetuning on WavLLM speech instruction data. SFT-LLM-LoRA is experimenting with the LoRA of rank 32 on Phi3.1-mini to align with the WavLLM [23] settings.
@@
For the ASR tasks, we use the repeat prompt during evaluation. The poor performance of E1 on Librispeech is due to the hallucinations in addition to the ASR transcriptions. An example of such hallucinations is shown in Box 1. In other words, the model cannot follow the exact prompt to repeat and tends to speak more. This behavior is typically observed in LLM. Such hallucinations are hard to quantify so that we don’t calculate IFR for the ASR tasks. Nevertheless, the ASR evaluations on audio-first training are still in a zero-shot manner and E1 should have a low IFR. For ST and SQA tasks, we can easily calculate the IFR using the method in Section III-C. We can observe from the table that AlignFormer can greatly improve the IFR for these tasks. For audio-first training, AlignFormer achieves a near 100%. For instruction-first training, AlignFormer improves the IFR to an average of 49%. We should specifically notice that the IFR on Gaokao QA has been improved to a non-zero IFR which is a game-changing improvement. This can be an important step towards the perfect speech and text modality matching in the LLM space and we show that it might be possible to achieve. Quality-wise, the performance of AlignFormer still encounters information loss problem because of lower token rate. For example, on CoVoST2 en-de evaluation, the BLEU scores are lower than the baselines even the IFR is higher. We still see the AlignFormer is promising because on MuST-C, the AlignFormer achieves a better BLEU than the cascades ASR + LLM system.
@@
There is an interesting finding on the SQA tasks. For the emotion tasks, there are four categories. The accuracy of zero-shot emotion evaluation shows that the model can infer the emotion from the ASR transcriptions to some degree (better than random guess with 25% for 4 categories). However, for the speaker verification task, the accuracy is similar to random guess (yes or no), indicating that the model is hard to infer speaker information from text. It provides us insights into the model development for emotion and speaker verification tasks.
When AlignFormer provides limited para-lingual information for LLM (e.g. random guess on the speaker verification task), a natural question is why not feeding the ASR output from CTC head to LLM for 100% IFR. In fact, the advantage of AlignFormer is the structure flexibility of learning over latent space instead of discrete token space with ASR outputs. The para-lingual information in the audio is able to be retrieved again from the latent space through task-specific training. To verify the hypothesis, we first evaluated AlignFormer (E1-audio-first) on WavLLM test data by feeding the CTC ASR output to LLM. Later, we conducted an additional supervised finetuning (SFT) stage with the WavLLM [23] instructional training data to show the advantage of AlignFormer operating in the latent space. During SFT, we freeze the speech encoder and CTC head when most of the SFT data provides no ASR supervision for learning CTC head. We either freeze LLM and tune the AlignFormer or add LoRA module to LLM to align with the setting in the WavLLM paper.
The results are presented in the last three columns of Table III. We observe two advantages of AlignFormer based on the results: (1) although CTC out + LLM offers good zero-shot capability for semantic heavy ST and SQA tasks, the performance on the ASR tasks with in-domain training data is much worse than the AlignFormer. The error propagation in cascades ASR + LLM solution can be reduced by learning over the latent space in AlignFormer. On the emotion and speaker verification tasks, CTC out + LLM intuitively performs bad, similarly to the AlignFormer. (2) When the AlignFormer is further trained with WavLLM SFT data, we can see that AlignFormer achieves better performance than WavLLM in most of the tasks. The gap on Emotion task is even reduced, indicating the emotion information is obtained through the adapter. However, the speaker information is still hard to be entirely retrieved from the top speech encoder layer to be on par with WavLLM structure that has dual encoders (semantic and acoustic encoders). This might be because the speech encoder is learned with ASR supervision in the alignment stage and is freezed during the SFT stage so it contains majorly semantic information. We believe the issue can be mitigated by incorporating some forms of speaker embeddings in AlignFormer design, which is left for future work.
Overall, we believe the advantage of AlignFormer to learn modality matching in the latent space. It can reduce the error propagation existing in ASR + LLM solution for in-domain training tasks. Additionally, the AlignFormer is flexible to finetune with instructional data to retrieve para-lingual information, while ASR +LLM solution cannot. Finally, using CTC text output as the input to LLM is a cascaded system while AlignFormer is a single model that has more potential to be optimized in an end-to-end way for more complicated speech tasks.